Two of the top grossing concert tours were by a jazz band and a rock band. Together the two visited 173 cities. The jazz band visited 93 more cities than the rock band.
how many cities did the jazz band visit?
how may cities did the rock band visit?

Answer by stanbon(75887)

Two of the top grossing concert tours were by a jazz band and a rock band. Together the two visited 173 cities. The jazz band visited 93 more cities than the rock band.
how many cities did the jazz band visit?
how may cities did the rock band visit?
------------
Equations:
r + j = 173
j = r + 93
-------------------
Substitute for "j" and solve for "r":
---
r + (r+93) = 173
2r = 80
r = 40 (# of rock band concerts)
---
j = r+93 = 40+93 = 133 (# of jazz band concerts)
================================================
Cheers,
Stan H.